OpenStudy (anonymous):

1/x-1 + 1/x+2 = 5/4 solve for all solutions?

OpenStudy (anonymous):

Multiply both sides by 4(x-1)(x+2) and see what you can get

OpenStudy (anonymous):

\[ 4 (x-1) (x+2) \left(\frac{1}{x+2}+\frac{1}{ x-1}\right)=\frac{5}{4} 4 (x-1) (x+2) \]

OpenStudy (anonymous):

\[ \frac{4 (x-1) (x+2)}{x-1}+\frac{4 (x-1) (x+2)}{x+2}=\frac{5}{4} 4 (x-1) (x+2) \] Simplify now.

OpenStudy (anonymous):

\[ 4 (x-1)+4 (x+2)=5 (x-1) (x+2) \]

OpenStudy (anonymous):

\[ 8 x+4=5 x^2+5 x-10 \]

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (anonymous):

yeah thx i understand now

hero (hero):

Actually x = 2 is one of the solutions since 2 + 2 = 4

OpenStudy (anonymous):

I thought x-2 was one of the denominator. I should have looked up.

hero (hero):

Yeah, I do the fractions on paper first before posting now to avoid mistakes like that. Actually, I have a technique that I use to make sure any extraneous roots are canceled before solving so I never run into that problem.

hero (hero):

And in this case, there were no extraneous roots.

OpenStudy (anonymous):

100%